OpenStudy (anonymous):

z varies directly with x and inversely with y. When x=4 and y=10, z=2. Find z when x= 5 and y=3.

OpenStudy (anonymous):

this means zy/x = constant

OpenStudy (anonymous):

hence zy/x , when z is 2, y=10 and x=4 is (2)(10)/4 = 5 now put zy/x = 5 at x=5, y=3 and find z

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

hold on im trying to understand :b

OpenStudy (anonymous):

i dont get the second part

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

now put zy/x = 5 at x=5, y=3 and find z <---- this

OpenStudy (anonymous):

see...we know that the expression zy/x is always equal to 5 dont we?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

soo... what do we do

OpenStudy (anonymous):

so zy/x = 5 means z(3)/5 =5 (plugging in x and y) now find z

OpenStudy (anonymous):

ohh okay. thank you

OpenStudy (anonymous):

not a problem :)

OpenStudy (nubeer):

lets suppose z=kx and z=u/y, k and u are constants ... when z=2,x=4,y=10 we get z=0.5x and z=20/y now just put the x=5 and y =3 u will get z=5/2 and z=20/3

OpenStudy (anonymous):

i dont think im getting the right answer :'(

OpenStudy (anonymous):

nvm igot it!

OpenStudy (nubeer):

ohh great if u got it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

for helping me

OpenStudy (anonymous):

wait soo i ended up with... z= 5 2/5 times 5 over 3 butt now idk how to do that on the calculator cause i know im doing it wrong...

OpenStudy (nubeer):

hmm i ddnt get what u asking can u plz explain lil more?